Reddick Inc. purchased machinery on 1/1/2018 for $220,000. Assuming no salvage value, JV depreciated the asset straight line over a 4-year life for Book purposes, and 3-year life for Tax purposes.  Assume a zero residual for both book and tax.  

Required:  (a) What would be the balance in the Deferred Tax Liability account at the end of each year, assuming a 30% tax rate?  (b) What would be the journal entry if Reddick sold the asset for $30,000 at the beginning of year 5 – be sure to consider the effect on taxes?
Answer #1
Answers:
(a) Calculation of yearly depreciation under straight line method:
Depreciation = (Acquisition value of the asset - Salvage value of the asset) / useful life of the asset
Machinery value = 220000
Salvage Value = 0
Life of the asset as per book = 4 years
Life of the asset as per tax = 3 years
Depreciation as per book = 220000/4 = 55000
Depreciation as per book = 220000/3 = 73333.33
Deferred tax liability balance at the end of each year as below:
Year Depreciation as per book Depreciation as per Tax Depreciation difference Tax rate Deferred tax liability Closing balance at the end of the each year
2018 55000 73333.33 -18333.33 30% -5500 -5500
2019 55000 73333.33 -18333.33 30% -5500 -11000
2020 55000 73333.33 -18333.33 30% -5500 -16500
2021 55000 55000 30% 16500 0
220000 220000
(b) Journal entry for Asset sold on profit:
Sl. No. General Ledger Debit Credit
1 Cash Account 30000
Machinery Account 30000
(Being asset sold)
2 Machinery Account 30000
Profit and Loss account 30000
(Being profit on the sale of the asset)
3 Income tax expenses 9000
Income tax provision 9000
(Being Tax on profit on sale of asset, 30000 X 30%)